OpenStudy (anonymous):

solve: lim->-∞ l3-xl + x

OpenStudy (anonymous):

Do you use piecewise funtion?

OpenStudy (anonymous):

Please explain step by step

OpenStudy (shubhamsrg):

try substituting x->-x because of which your limit becomes ∞ from -∞

OpenStudy (anonymous):

i don't get it, why would i do this?

OpenStudy (shubhamsrg):

then you'll have |3+x| -x with x->inf this'll help you take care of the mod

OpenStudy (anonymous):

i've already solved for +inf, answer is inf.

OpenStudy (shubhamsrg):

answer according to be should not be inf

terenzreignz (terenzreignz):

I thought we'd already done this? http://openstudy.com/study#/updates/51c6fb73e4b069eb00c5dc53

OpenStudy (anonymous):

yes, but i'm still struggling

OpenStudy (anonymous):

i'm just more confused now

terenzreignz (terenzreignz):

Okay, so let's have a look at that function again... \[\large f(x) =|3-x|+x\]

terenzreignz (terenzreignz):

And see how it behaves. Int particular, the |3 - x| bit.

OpenStudy (anonymous):

alright; so what do i do to solve it?

OpenStudy (anonymous):

are there different methods?

terenzreignz (terenzreignz):

I'm sure, but I'm not familiar with them. Better check back with @shubhamsrg :D

OpenStudy (shubhamsrg):

i just had 1 method in mind, to substitute x for -x :|

terenzreignz (terenzreignz):

still different :) I expressed as... \[\Large f(x) = |3-x|+x = \left\{\begin{matrix}3& \qquad x\le3\\2x-3&\qquad x>3\end{matrix}\right.\]

OpenStudy (shubhamsrg):

(Y)

OpenStudy (anonymous):

shouldn't it just be: l3-xl = (3-x) if x >3 -(3-x) if x<3

terenzreignz (terenzreignz):

Okay, here's the thing... first things first, |a| = a if a>0 |a| = -a if a<0 catch me so far?

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

Okay... now if x > 3 Then... 0 > 3-x correct?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well then, flipping it around, we get 3 - x < 0

OpenStudy (anonymous):

alright

terenzreignz (terenzreignz):

Well then, 3 - x < 0 which means |3 - x| = -(3-x) = x - 3

OpenStudy (anonymous):

wow what

OpenStudy (anonymous):

oh ok, i get it

terenzreignz (terenzreignz):

We said that if a < 0 then |a| = -a

terenzreignz (terenzreignz):

3-x is no exception. Since 3 - x < 0 then |3 - x| = -(3 - x) = x - 3

OpenStudy (anonymous):

yes, you put the minus to get a positive again because its an absolute value

terenzreignz (terenzreignz):

Yup. So you understand that IF x > 3 then |3 - x| = x - 3 right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Okay, from that same logic, what if x < 3?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

i thought -3-x

terenzreignz (terenzreignz):

well, let's subtract x from both sides giving us 0 < 3 - x Or 3 - x > 0 From this, what can we conclude?

OpenStudy (anonymous):

i give up

terenzreignz (terenzreignz):

a > 0 then |a| = ?

terenzreignz (terenzreignz):

And giving up is not allowed. Particularly during quizzes or exams :P

OpenStudy (anonymous):

just a

terenzreignz (terenzreignz):

Yup. so |3 - x| = ?

OpenStudy (anonymous):

3-x?

terenzreignz (terenzreignz):

Yes. So, recap so far. If x > 3 then |3 - x| = x - 3 And if x < 3 then |3 - x| = 3 - x catch me so far?

OpenStudy (anonymous):

they're both the same?

OpenStudy (anonymous):

i should write this down, might make more sense

terenzreignz (terenzreignz):

No, they're not. They're negatives of each other.

OpenStudy (anonymous):

how come you say: -(3-x) = x-3

terenzreignz (terenzreignz):

What else could it be? \[\Large \color{red}-(3-x)=\color{red}-3 \color{red}--x = -3+x = x -3\]

OpenStudy (anonymous):

haha, stupid question (tired)

OpenStudy (anonymous):

but anyway, i get it

OpenStudy (anonymous):

i wrote it down and it makes sense

terenzreignz (terenzreignz):

Okay, that said... If x > 3 implies |3 - x| = x - 3 then what bodes for |3 - x| + x = ? (PS avoid maths when you're tired. It could lead to so many errors.) Maybe you should rest? :D

OpenStudy (anonymous):

haha, no i need to get this ;)

terenzreignz (terenzreignz):

Well then, I had a question for you ^

OpenStudy (anonymous):

2x - 3

terenzreignz (terenzreignz):

How?

OpenStudy (anonymous):

x - 3 + x

terenzreignz (terenzreignz):

Very good. I just wanted to make sure you understand :) And what does x < 3 imply for |3 - x| + x ?

OpenStudy (anonymous):

3 - x + x = 3

terenzreignz (terenzreignz):

Good :) So now, I take it you understand why \[\Large f(x) = |3-x|+x= \left\{\begin{matrix}3& \qquad x\le3\\2x-3&\qquad x>3\end{matrix}\right.\]?

OpenStudy (anonymous):

phew, yes

terenzreignz (terenzreignz):

Well then, you'll notice that as x goes to negative infinity, ALL THOSE NUMBERS are going to be less than 3, so the value for f(x) will always be...?

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

And hence the limit :)

OpenStudy (anonymous):

aaalright, thx!

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

thx for your patience too

terenzreignz (terenzreignz):

Sure :)

OpenStudy (anonymous):

and the question can be closed! woohoo

terenzreignz (terenzreignz):

I'd just like to point out that when x = 3 (we didn't cover this) then |3 - x| + x = 3 = 2x - 3 That is to say, when x = 3 it doesn't matter if you define f(x) to be 3 or 2x - 3 it'll be the same value.

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

thx again

terenzreignz (terenzreignz):

But that doesn't affect the nature of this question. It was just a thought.

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

Close away :)

OpenStudy (anonymous):

alright, till next time!